find a quadratic equation with roots -1-4i and -1+4i please help!
Answer by stanbon(75887)

find a quadratic equation with roots -1-4i and -1+4i
------
f(x) = (x-(-1-4i))(x-(-1+4i))
-----
= ((x+1)+4i)((x+1)-4i)
-----
= (x+1)^2 - (4i)^2
----
= (x+1)^2 + 16
----
= x^2 + 2x + 17
==================
Cheers,
Stan H.
==================